# CaTS-Bench: Can Language Models Describe Numeric Time Series?

### v1

@@ Abstract
Time series captioning, the task of describing numeric time series in natural language, requires numeric reasoning, trend interpretation, and contextual understanding. Existing benchmarks, however, often rely on synthetic data or overly simplistic captions, and typically neglect metadata and visual representations. To close this gap, we introduce CaTS-Bench, the first large-scale, real-world benchmark for Context-aware Time Series captioning. CaTS-Bench is derived from $11$ diverse datasets reframed as captioning and Q&A tasks, comprising roughly $465k$ training and $105k$ test timestamps. Each sample includes a numeric series segment, contextual metadata, a line-chart image, and a caption. A key contribution of this work is the scalable pipeline used to generate reference captions: while most references are produced by an oracle LLM and verified through factual checks, human indistinguishability studies, and diversity analyses, we also provide a human-revisited subset of $579$ test captions, refined from LLM outputs to ensure accuracy and human-like style. Beyond captioning, CaTS-Bench offers $460$ multiple-choice questions targeting deeper aspects of time series reasoning. We further propose new tailored evaluation metrics and benchmark leading VLMs, highlighting both their strengths and persistent limitations. Together, these contributions establish CaTS-Bench and its captioning pipeline as a reliable and extensible foundation for future research at the intersection of time series analysis and foundation models.
Figure 1: Overview of CaTS-Bench. It features diverse domains, provides training and benchmark data, and formulates five challenging tasks, with time series captioning as the primary one.
@@ 1 Introduction
Effective interpretation of time series data is a cornerstone of decision‑making in domains ranging from financial markets and healthcare monitoring to climate analysis and industrial automation. Yet, distilling raw numeric sequences into concise, human‑readable summaries remains a labor‑intensive task, requiring domain expertise, statistical know‑how, and careful visualization. Automating this process through time series captioning (TSC) not only accelerates insight discovery but also democratizes access to complex temporal analytics, enabling non‑experts to ask natural‑language questions and receive meaningful explanations without writing code or inspecting raw charts.
Large language models (LLMs) and vision‑language models (VLMs) have demonstrated remarkable prowess in text generation and visual reasoning, respectively. However, when applied to time series, they reveal critical deficiencies: LLMs exhibit well-documented limitations in precise numeric extrapolation, temporal continuity, and uncertainty quantification (Tang et al., 2025; Merrill et al., 2024; Tan et al., 2024; Cao & Wang, 2024). While VLMs have shown promise in visual pattern recognition tasks such as trend and anomaly detection from plots (Zhou & Yu, 2025), their capacity for fine-grained numeric time series reasoning remains largely underexplored. These limitations underscore a broader challenge: existing evaluation resources fail to reflect the complexity of real‑world temporal signals, leaving model improvements unguided by the demands of true data‑driven applications.
In response, the community has proposed Time Series Captioning (TSC) as a more natural task for foundation models, leveraging their generative and reasoning capabilities to narrate trends, anomalies, and context in prose (Trabelsi et al., 2025; Jhamtani & Berg-Kirkpatrick, 2021). However, current benchmarks remain narrow, often synthetic or restricted to simple trend labels, and exclude rich metadata or visual modalities. Consequently, progress in model architecture, pretraining, or finetuning cannot be measured against challenges that mirror real deployment scenarios, slowing adoption in high‑stakes sectors where accurate temporal interpretation is essential.
To fill this gap, we introduce CaTS-Bench, the first large-scale, multimodal benchmark explicitly designed for context-aware time series captioning and reasoning. We define “context-aware” to mean that captions are informed by both the metadata (units, domain labels, dates, region, etc.) and visual cues that provide semantic and numeric grounding. By mining $11$ real-world datasets across various domains, CaTS-Bench provides $20k$ triplet samples drawn from $570k$ time steps of curated data, each paired with (1) rich metadata containing contextual information, units, and domain-specific cues (Dong et al., 2024; Wang et al., 2024); (2) a corresponding line plot image, enabling the use of VLMs (Chen et al., 2024a; Zhou & Yu, 2025); and (3) a reference caption produced by a scalable oracle-based pipeline and validated through factual checks, human indistinguishability studies, and diversity analyses. To further strengthen reliability, we additionally release a human-revisited subset of test captions: sampled from multiple LLM candidates and carefully edited by the authors to remove inaccuracies, speculative claims, and linguistic repetitions. This subset complements the larger benchmark with high-fidelity, human-styled references. Beyond captioning, CaTS-Bench also includes $460$ challenging multiple-choice questions spanning time series matching, caption matching, plot matching, and comparative reasoning, designed to expose models’ blind spots in numeric precision and multimodal alignment. All data samples are made available here.
We further propose new evaluation metrics tailored to time series captioning that move past generic N-gram overlap to reward numeric fidelity and coverage. Our comprehensive experiments on leading VLMs reveal that, in both zero-shot and finetuned settings, models can produce fluent text but fail to reliably capture quantitative details without specialized adaptation. A key finding is that VLMs fail to effectively leverage the visual cues provided for time series captioning, pointing to a significant limitation in current multimodal architectures. Our analysis identifies clear room for improvement, such as better leveraging visual cues, enhancing multimodal alignment, and incorporating dedicated numeric reasoning modules. These findings pave the way for a new generation of foundation models capable of translating complex temporal data into actionable narratives.
In summary, the contributions of this paper are:
@@
    Scalable Captioning Pipeline: A reproducible pipeline for generating high-quality time series captions. It anchors LLM outputs in factual metadata, validates them through factual checks, human indistinguishability studies, and diversity analyses, and is extensible to new datasets.
@@
    CaTS-Bench: A multimodal, context-aware benchmark for time series captioning and reasoning, featuring time series segments, rich metadata, visual plots, and factually grounded captions. Most references are LLM-generated via the pipeline, while a curated subset of human-revisited test captions ensures high-fidelity, human-styled references alongside the larger benchmark.
@@
    Diagnostic Q&A Suite: Four multiple-choice tasks designed to isolate capabilities in series matching, caption grounding, visual reasoning, and comparative analysis.
  4. 4.
    Comprehensive Evaluation: Zero-shot and finetuned assessments of state-of-the-art VLMs, revealing strengths, failure modes, and clear directions to advance time series understanding.
@@ 2 Related Work
LLMs are increasingly being repurposed for time series analysis (Zhang et al., 2024; Liu et al., 2024a), with early efforts primarily focused on forecasting. These approaches span prompt engineering (Liu et al., 2024a; Chatzigeorgakidis et al., 2024), modality alignment (Liu et al., 2024b; Sun et al., 2023; Liu et al., 2024c; Pan et al., 2024), discretization (Ansari et al., 2024; Jin et al., 2024), and specialized finetuning (Zhou et al., 2023; Chang et al., 2023). Such studies highlight that LLMs pretrained on text can reason over temporal data, but subsequent work also shows consistent weaknesses in handling long-range dependencies, numeric precision, and structured reasoning, particularly in forecasting and anomaly detection (Tang et al., 2025; Merrill et al., 2024; Tan et al., 2024; Cao & Wang, 2024; Zeng et al., 2023).
@@
Dataset # Timesteps Modality Sources Metadata Captions TSC Q&A
TADACap (Fons et al., 2024) N/A Visual 4 Minimal Patterns Only
TRUCE (Jhamtani & Berg-Kirkpatrick, 2021) 34k Numeric 2 Patterns Only
TACO (Dohi et al., 2025) 2.46b Numeric 8 Expressive
CaTS-Bench 570k Numeric + Text + Visual 11 Rich Expressive
Building on these foundations, researchers have explored Time Series Captioning (TSC), a task more aligned with the generative strengths of language models. TSLM (Trabelsi et al., 2025) introduces an encoder–decoder trained on synthetic cross-modal data; TADACap (Fons et al., 2024) retrieves domain-aware captions for visualized time series; TRUCE (Jhamtani & Berg-Kirkpatrick, 2021) employs a truth-conditional framework to validate simple trend patterns; and TACO (Dohi et al., 2025) scales up caption corpora using LLM-based synthetic generation. While each provides valuable insights, they remain limited in scope: TADACap and TRUCE are domain-specific and pattern-oriented, while TACO’s reliance on templates restricts contextual richness (See Table 1).
Beyond these, standard time-series archives such as UCR (Chen et al., 2015), UEA (Bagnall et al., 2018), and Monash (Godahewa et al., 2021) support classification and forecasting but not generative captioning. Similarly, benchmarks like PISA (Xue & Salim, 2023) target prompt-based forecasting, omitting metadata entirely. Recent evidence shows that incorporating auxiliary modalities (metadata, domain context, or visual renderings) can significantly improve both interpretability and predictive performance (Zhou & Yu, 2025; Dong et al., 2024; Chen et al., 2024a; Wang et al., 2024; Kim et al., 2024; Williams et al., 2024; Liu et al., 2025; Tang et al., 2023). Yet no benchmark to date integrates large-scale numeric series, expressive captions, rich metadata, and multimodal grounding. CaTS-Bench fills this gap by offering the first benchmark that unifies numeric time series, metadata, and visuals with both expressive captions and Q&A tasks for systematic evaluation for TSC.
@@ 3 CaTS-Bench
In this section, we illustrate the entire data curation pipeline and the design of the benchmark tasks. While examples generated from this pipeline can be directly used for TSC evaluation, we further enrich the scope of CaTS-Bench by providing an additional suite of Q&A tasks constructed from the same data, enabling a more fine-grained examination of time series and caption reasoning abilities.
3.1 Data Curation
We build CaTS-Bench, a comprehensive benchmark curated from $11$ diverse real-world source datasets spanning domains: climate (Jha, 2023; Ritchie, 2021), safety (of Los Angeles, n.d.; of Public Health, n.d.), USA border crossing (U.S. Department of Transportation, n.d.), demography (Aziz, 1985), health (European Centre for Disease Prevention and Control, 2024; Food and Agriculture Organization of the United Nations, 2024), sales (Hassan, 2020; Chen, 2015), and agriculture (USDA Economic Research Service, 2024). See Appendix B for more details on the source datasets. The overall data pipeline is shown in Figure 2. Each source dataset provides a full-length time series per entity (e.g., country, city, product), and to generate samples, we apply a random window cropping strategy. For each dataset, we define a valid range of window lengths and randomly select a size for each crop; see Appendix C for our range calculation. The number of windows sampled from a dataset depends on its total time steps, ensuring fair representation. The domain-specific number and lengths of the time series windows are illustrated in Table 2. Each time series window is augmented with a metadata JSON file with contextual information (domain, location, start time, etc.), a line plot image with randomized visual style (color, width, figure size), a ground truth caption produced by querying an oracle LLM (Gemini 2.0 Flash) with a structured prompt that includes: (i) the serialized numeric values of the cropped segment and (ii) metadata enriched with numericly grounded information, including both the historical and sample-specific mean, standard deviation, minimum, and maximum. An example of the prompt is available in Appendix N.1.
Figure 2: Overview of the CaTS-Bench semi-synthetic data generation pipeline. A time series window is cropped, metadata is attached, and an oracle LLM generates a reference caption. See Appendix L for examples and Appendix H for the quality verification protocol.
We emphasize that time series captioning lacks inherent ground truth at the level of a single canonical description: multiple valid ways exist to describe the same series depending on focus and phrasing. To provide consistent references at scale, our primary captions are generated by an oracle LLM, but anchored strictly in the underlying data. The oracle receives full contextual metadata (not available at evaluation time) and is instructed not to include any external knowledge, ensuring captions remain factual and context-grounded. This design makes captions a practical proxy for evaluation and challenges models to reason from multimodal inputs rather than mimic the oracle. Furthermore, we randomize time series window sizes and plot styles to prevent overfitting and better reflect real-world variability in length and visualization styles.

### v2

# CaTS-Bench: Can Language Models Describe Time Series?

@@ Abstract
Time series captioning, the task of describing time series in natural language, requires numeric and temporal reasoning, trend interpretation, and contextual understanding. Existing benchmarks, however, often rely on fully synthetic or generic captions, and typically neglect metadata and visual representations. We introduce CaTS-Bench, a comprehensive benchmark for Context-aware Time Series reasoning across $11$ diverse domains, centered on a gold-standard evaluation set of $1746$ human-rewritten captions that measure how effectively models translate numeric trends into immediately interpretable narratives. To address the scarcity of human-annotated data, we also propose a scalable pipeline for generating high-fidelity synthetic captions, the quality of which we validate. We evaluate leading Vision-Language Models on our benchmark, revealing that even proprietary models struggle to capture numeric nuances in temporal descriptions, while finetuning open-source models on synthetic data yields substantial performance gains. Finally, release a diagnostic suite of $910$ multiple-choice questions and tailored numeric metrics to gauge time-series-specific reasoning capabilities, establishing CaTS-Bench as a reliable foundation for grounded, multimodal language generation in numeric domains.
CaTS-Bench: Can Language Models Describe Time Series?
Luca Zhou*1, Pratham Yashwante*2, Marshall Fisher2, Alessio Sampieri3, Zihao Zhou2, Fabio Galasso1, Rose Yu2 1Sapienza University of Rome, 2UC San Diego, 3ItalAI Labs Correspondence: luca.zhou@uniroma1.it
Figure 1: Overview of CaTS-Bench. It features $11$ domains organized into 77 macro domains, provides human-rewritten and synthetic data, and formulates six challenging tasks, with time series captioning as the primary one.
@@ 1 Introduction
Effective interpretation of time series data underpins decision-making in domains such as finance, healthcare, and climate analysis. Although raw numeric sequences are abundant, turning them into concise, human-readable summaries is labor-intensive and demands domain expertise and statistical literacy. In practice, analysts, policymakers, journalists, and practitioners rarely inspect raw sequences directly, instead relying on textual descriptions to quickly assess trends, anomalies, and changes before deciding whether deeper analysis is needed. Automating this process through time series captioning (TSC) not only accelerates insight discovery but also democratizes access to complex temporal analytics, which enables non‑experts to ask natural‑language questions and receive explanations without writing code or inspecting raw charts.
Recent advances in Large Language Models (LLMs) and Vision-Language Models (VLMs) have transformed grounded language generation, yet these models show critical weaknesses in precise reasoning and factual grounding on numeric time series (tang2025time; cao2024evaluation). Progress in TSC is bottlenecked by the lack of high-quality, linguistically diverse datasets and benchmarks: existing resources are largely synthetic, template-based, or omit contextual metadata (e.g., units, domain-specific labels) that humans rely on, leading models to produce generic captions that miss the underlying narrative.
To address this, we introduce CaTS-Bench, a context-rich multimodal benchmark for context-aware time series captioning. Here, “context-aware” denotes synthesizing information from numeric series, visual plots, and contextual metadata. Rather than assuming that access to multiple modalities implies effective multimodal reasoning, we explicitly probe whether recent VLMs properly digest these signals. Unlike previous efforts that rely solely on simplistic or synthetic data, our work centers on a new gold-standard evaluation set of $1746$ human-rewritten (HR) captions. These captions, curated across $11$ real-world domains, provide a rigorous baseline to measure linguistic richness, factual and numeric precision, and human-like reasoning in temporal descriptions.
Recognizing the prohibitive cost of large-scale human annotation, a second key contribution of this work is a scalable data augmentation pipeline that uses an oracle LLM, grounded in factual metadata and validated via human verification studies, to generate high-fidelity synthetic training captions. Experiments show that zero-shot VLMs struggle with the complexity of HR descriptions, whereas finetuning on our synthetic data yields substantial gains, particularly in semantic and numeric metrics on the HR benchmark, positioning the pipeline as an effective solution to data scarcity for TSC.
Beyond captioning, CaTS-Bench provides a diagnostic suite of $910$ multiple-choice questions (MCQs) targeting numeric precision and multimodal alignment, together with tailored evaluation metrics that prioritize numeric fidelity. Our results reveal that current VLMs often fail to exploit visual cues for TSC and struggle even on simple diagnostic MCQ tasks, such as correctly mapping values to time stamps and comparing basic statistical properties, highlighting persistent gaps in grounded multimodal reasoning. Overall, our contributions are:
@@
    CaTS-Bench: A multimodal, 11-domain benchmark featuring a 1746-sample human-rewritten evaluation set for TSC and $910$ MCQs designed to rigorously measure linguistically rich and factually grounded temporal reasoning.
@@
    Scalable Augmentation Pipeline: A framework for generating high-fidelity synthetic data. We empirically demonstrate that finetuning on this synthetic data significantly boosts model performance on our human-rewritten test set.
@@
    Diagnostic Analysis & Metrics: A suite of five diagnostic Q&A tasks and tailored numeric fidelity metrics that expose systematic failures in visual grounding and temporal alignment, providing a rigorous diagnostic baseline for grounded multimodal free-form generation.
@@ 2 Related Work
Dataset Data Nature Modalities Domains Context Captions Q&A
TaxoSynth (fons2024evaluating) Synthetic Numeric + Text N/A Minimal Simplistic
TADACap (fons2024tadacap) Partially Synthetic Visual 4 Minimal Simplistic
TRUCE (jhamtani2021truth) Partially Synthetic Numeric 2 Simplistic
TACO (dohi2025domain) Mostly Synthetic Numeric 8 Expressive
BEDTime sen2025bedtimeunifiedbenchmarkautomatically Partially Synthetic Numeric + Text 6 Simplistic
CaTS-Bench (Ours) HR + Synthetic Numeric + Text + Visual 11 Rich Expressive
LLMs for Time Series Analysis. LLMs are increasingly repurposed for temporal reasoning, with early efforts focused on forecasting and anomaly detection (zhang2024large; liu2024large). Methodological approaches include specialized prompt engineering (liu2024large; chatzigeorgakidis2024multicast), modality alignment (liu2024timecma; suntest; liu2024taming; pan2024s; cai2024jolt), numeric discretization (ansari2024chronos; jin2023time), and parameter-efficient finetuning (zhou2023one; chang2023llm4ts). While these studies demonstrate that models pretrained on massive text corpora can inherit temporal reasoning capabilities, they also highlight persistent "hallucinations" in numeric extrapolation and difficulty handling long-range dependencies (tang2025time; merrill2024language; tan2024are; cao2024evaluation; zeng2023transformers). We shift the focus from purely predictive tasks to generative description, exploring how these models translate numeric signals into language.
Related Benchmarks and TSC. Traditional time-series archives (UCRArchive; bagnall2018uea; godahewa2021monash) focus on discriminative tasks, while prompt-based forecasting benchmarks like PISA (xue2023promptcast) lack the linguistic pairings and metadata required for descriptive tasks. While recent TSC-specific frameworks have pioneered the area (fons2024tadacap; jhamtani2021truth; fons2024evaluating; dohi2025domain; sen2025bedtimeunifiedbenchmarkautomatically), they typically rely on simplistic trend patterns or template-based synthetic generation, failing to capture the descriptive nuance of human discourse (see Table 1). Furthermore, although integrating auxiliary modalities, such as metadata and visual renderings, is vital for grounded interpretability (zhou2025can; dong2024metadata; chen2024visionts; wang2024news; kim2024multi; williams2024contextkeybenchmarkforecasting; liu2025can; tang2023vistext), a unified benchmark integrating numeric data, expressive human captions, and multimodal grounding has remained elusive. CaTS-Bench addresses this gap with a unified benchmark for grounded TSC that integrates numeric data, visual representations, contextual metadata, and human language.
@@ 3 CaTS-Bench
Here, we describe the construction of CaTS-Bench, starting from raw numeric series to our dual-stream captioning framework. CaTS-Bench provides a HR gold standard for evaluation, alongside a large-scale synthetic corpus for data augmentation. Moreover, we further enrich the scope of CaTS-Bench by providing an additional suite of Q&A tasks constructed from the same data, enabling a more fine-grained examination of time series and caption reasoning abilities.
3.1 Base Data Curation
We work with $11$ diverse real-world source datasets that span climate, public safety, border crossings, demography, health, sales, and agriculture. See Appendix B for details of the source datasets. This variety ensures that models must generalize across different temporal scales, units, and domains.
The curation process begins by generating base data triplets. For each entity in a source dataset (e.g., a specific country or product), we apply a random window cropping strategy to extract segments of varying lengths. For each segment, we construct a triplet consisting of (i) the raw numeric time series, (ii) a structured metadata JSON containing domain-specific context (e.g., units, location, temporal bounds) and statistical primitives (mean, dispersion, minimum, and maximum), and (iii) a line plot rendering of the series with randomized visual styles (color, width, and size) to promote robustness to visual variation. The overview is shown in 2. To prevent information leakage, we partition the data temporally: the first 80% of each series is reserved for training/augmentation, while the final 20% is used for the actual benchmark data. The sizes of data splits are reported in Table 2.